OpenStudy (anonymous):

ok i am completely throw for a loop with this one please help Let f(x) = 12 over the quantity of 4 x + 2 Find f(-1).

hartnn (hartnn):

to get f(-1) , just plug in x=-1 in f(x) what do u get ?

OpenStudy (anonymous):

it is like a fraction im so confused

hartnn (hartnn):

then lets take the denominator only! 4x+2 put x=-1 what do u get ?

OpenStudy (anonymous):

3+2

hartnn (hartnn):

4(-1) means 4 is multiplied by -1 so, 4*(-1) =... ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

no -3 i mean

hartnn (hartnn):

no... 4-1 =3 but not 4*(-1)

hartnn (hartnn):

what is 4*1 = ... ?

OpenStudy (anonymous):

4

hartnn (hartnn):

yes, so 4*(-1) = - 4*1 = ... ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so it should b -4

hartnn (hartnn):

yes

hartnn (hartnn):

so, 4x+2 = -4+2 = ... ?

OpenStudy (anonymous):

-2

hartnn (hartnn):

yes, so f(-1) = 12/(-2) = -12/2 = ... ?

OpenStudy (anonymous):

oh ok so i should just like move the minus sign

OpenStudy (anonymous):

-6

hartnn (hartnn):

thats correct! f(-1) = -6 :)

OpenStudy (anonymous):

ok i got it now tyvm

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

can u help me with one more please

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x).

OpenStudy (anonymous):

x^2 and the f^-1

OpenStudy (anonymous):

i completely forgot how to do exsponents

hartnn (hartnn):

so we take y= x^2-81 interchange 'x' and 'y' to get x = y^2 -81 now can you try to isolate y from this ?

OpenStudy (anonymous):

good god my brain hurts lol

hartnn (hartnn):

start by adding 81 on both sides

OpenStudy (anonymous):

so it would look like 81x=y^2-81

hartnn (hartnn):

no...

hartnn (hartnn):

you know what is "adding" ?

OpenStudy (anonymous):

yes i know how to add

hartnn (hartnn):

ok, so adding 81 to x will give you x+81 right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would b 81x

hartnn (hartnn):

no...81x means multiplying x with 81

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so just x+81

hartnn (hartnn):

x+81 = y^2 take square root on both sides

OpenStudy (anonymous):

how do i get into sq roots

OpenStudy (anonymous):

the only number that i know that relates to 81 is 9

hartnn (hartnn):

\(y^2= x+81 \\ y=\sqrt{x+81}\) thats it!

OpenStudy (anonymous):

ok stupider question how did u make sq root sign on puter

hartnn (hartnn):

here ? i wrote a code `\(\sqrt{...}\)` \(\sqrt{...}\)

OpenStudy (anonymous):

in my answer list this is there is this rt ±square root of the quantity x plus 81

hartnn (hartnn):

yes, when you take square root, you need to consider both + and -

OpenStudy (anonymous):

or is it one of theses 9square root of x 1 over quantity x squared minus 81 x squared over 81

hartnn (hartnn):

its this \(\pm \sqrt {x+81}\)

OpenStudy (anonymous):

i gonna have to get my kid to figure out how to put that code in on my laptop

OpenStudy (anonymous):

ok yes thats it exactly the way it looks on the answer thingy tysssvm

hartnn (hartnn):

welcome :)